OpenStudy (anonymous):

What is the tenth term in the sequence below? √3, 3, 3√3, . . . A. 1/243 B. 1/81 C. 243 D. 729

OpenStudy (anonymous):

it's a geometric sequence with common ratio = sqrt(3)

OpenStudy (anonymous):

But how do I find the answer

OpenStudy (anonymous):

a(n) = a1 * r^(n-1) you know a1,r and n, plug them in to find a(n)

OpenStudy (anonymous):

I'm sorry I'm not trying to make you give me the answer but Im honestly confused. What are the a1, r, and n?

OpenStudy (anonymous):

a1 = first term r = common ratio n = position of each term so a(10) = sqrt(3) (sqrt(3))^(10-1) = 243

OpenStudy (anonymous):

Omg thank you so much